### v2

@@ 3 Coderivatives of the second-order cone
$\partial_{B}\Pi_{{\mathcal{Q}}}({\boldsymbol{0}})=\{I_{1+r},O_{1+r}\}\cup\left\{\frac{1}{2}\begin{pmatrix}1&{\boldsymbol{w}}^{\top}\\ {\boldsymbol{w}}&2\alpha I_{r}+(1-2\alpha){\boldsymbol{w}}{\boldsymbol{w}}^{\top}\end{pmatrix}\,\middle|\,\begin{array}[]{ll}{\boldsymbol{w}}\in{\mathds{R}}^{r},\|{\boldsymbol{w}}\|=1,\\ \alpha\in[0,1]\end{array}\right\}.$ (3.1)
@@
$A({\boldsymbol{z}}):=I_{1+r}-\frac{1}{2}\begin{pmatrix}-1\\ \frac{\bar{{\boldsymbol{z}}}}{\|\bar{{\boldsymbol{z}}}\|}\end{pmatrix}\begin{pmatrix}-1&\frac{\bar{{\boldsymbol{z}}}^{\top}}{\|\bar{{\boldsymbol{z}}}\|}\end{pmatrix}\quad\mbox{and}\quad B({\boldsymbol{z}}):=\frac{1}{2}\begin{pmatrix}1\\ \frac{\bar{{\boldsymbol{z}}}}{\|\bar{{\boldsymbol{z}}}\|}\end{pmatrix}\begin{pmatrix}1&\frac{\bar{{\boldsymbol{z}}}^{\top}}{\|\bar{{\boldsymbol{z}}}\|}\end{pmatrix}.$ (3.2)
@@
${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})=\{{\boldsymbol{p}}\mid{\boldsymbol{p}}=-{\boldsymbol{q}}-{\boldsymbol{r}},-W{\boldsymbol{r}}={\boldsymbol{q}}\}=\{{\boldsymbol{p}}\mid W{\boldsymbol{p}}=(I-W){\boldsymbol{q}}\},\quad\forall\,{\boldsymbol{q}}\in{\mathds{R}}^{1+r}.$ (3.3)
@@
which, together with (3.3), implies the explicit formulation of ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})$.
@@
$-{\boldsymbol{q}}\in\begin{cases}{\sf conv}\{-{\boldsymbol{p}}-{\boldsymbol{q}},A({\boldsymbol{u}})(-{\boldsymbol{p}}-{\boldsymbol{q}})\},&\mbox{if }\langle(-{\boldsymbol{p}}-{\boldsymbol{q}}),(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|})\rangle\geq 0,\\ \{-{\boldsymbol{p}}-{\boldsymbol{q}},A({\boldsymbol{u}})(-{\boldsymbol{p}}-{\boldsymbol{q}})\},&\mbox{ otherwise},\end{cases}\quad\forall\,{\boldsymbol{p}},{\boldsymbol{q}}\in{\mathds{R}}^{1+r},$ (3.4)
@@
where the definition of $A({\boldsymbol{u}})$ comes from (3.2). Note that (3.4) holds if and only if
@@
${\boldsymbol{p}}=\frac{\varrho}{2}\begin{pmatrix}1\\ -\frac{\bar{{\boldsymbol{u}}}}{\|\bar{{\boldsymbol{u}}}\|}\end{pmatrix}\begin{pmatrix}1,-\frac{\bar{{\boldsymbol{u}}}^{\top}}{\|\bar{{\boldsymbol{u}}}\|}\end{pmatrix}{({\boldsymbol{p}}+{\boldsymbol{q}})},\ \varrho\in\begin{cases}[0,1],&\mbox{if }-\langle{\boldsymbol{p}},(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|})\rangle\geq\langle{\boldsymbol{q}},(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|})\rangle,\\ \{0,1\},&\mbox{ otherwise}.\end{cases}$ (3.5)
Moreover, (3.5) implies
$\begin{array}[]{ll}(1-\varrho)\langle(1;-\frac{\bar{{\boldsymbol{u}}}}{\|\bar{{\boldsymbol{u}}}\|}),{\boldsymbol{p}}\rangle=\varrho\langle(1;-\frac{\bar{{\boldsymbol{u}}}}{\|\bar{{\boldsymbol{u}}}\|}),{{\boldsymbol{q}}}\rangle.\end{array}$ (3.6)
@@
For a given vector ${\boldsymbol{q}}\in{\mathds{R}}^{1+r}$, we consider the following three cases, based on the value of $\langle(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}),{\boldsymbol{q}}\rangle$, to get the equivalent conditions of ${\boldsymbol{p}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})$ from (3.5):
@@
    If $\langle(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}),{\boldsymbol{q}}\rangle>0$, one can not take $\varrho=1$, or else (3.6) fails to hold. When $\varrho\in(0,1)$, by (3.6) one should have $\langle(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}),{\boldsymbol{p}}\rangle>0$, but one can not find a vector ${\boldsymbol{p}}$ such that (3.5) holds in this case. Therefore, one can only take $\varrho=0$, so that ${\boldsymbol{p}}={\boldsymbol{0}}$ by the equality in (3.5). In this case, it is easy to see that (3.5) holds. Thus, ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})=\{{\boldsymbol{0}}\}$.
@@
    If $\langle(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}),{\boldsymbol{q}}\rangle=0$, by (3.6) one has $\langle(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}),{\boldsymbol{p}}\rangle=0$ whenever $\varrho\in[0,1)$. In this case, (3.5) holds if and only if ${\boldsymbol{p}}={\boldsymbol{0}}$. Moreover, when $\varrho=1$, (3.6) always holds. In this case, (3.5) holds if and only if ${\boldsymbol{p}}=\frac{1}{2}(1,-\frac{\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|}){\boldsymbol{p}}(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|})$. Thus, ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})=\{\tau(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|})\mid\tau\in{\mathds{R}}\}$.
@@
    If $\langle(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}),{\boldsymbol{q}}\rangle<0$, one can not take $\varrho=1$, or else (3.6) fails to hold. When $\varrho=0$, it is easy to see that (3.5) holds if and only if ${\boldsymbol{p}}={\boldsymbol{0}}$. When $\varrho\in(0,1)$, by (3.6) one should have $\langle(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}),{\boldsymbol{p}}\rangle<0$. Therefore, (3.5) holds if and only if ${\boldsymbol{p}}=\tau(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|})$ with $\tau:=\frac{\varrho}{2}(1,-\frac{\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|})({\boldsymbol{p}}+{\boldsymbol{q}})\leq 0$. Note that such a condition is valid for any $\tau<0$ since $\varrho=\frac{2\tau}{2\tau+\langle(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}),{\boldsymbol{q}}\rangle}\in(0,1)$ always holds. Thus, ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})=\{\tau(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|})\mid\tau\leq 0\}$.
@@
$-{\boldsymbol{q}}\in\begin{cases}{\sf conv}\{{\boldsymbol{0}},B({\boldsymbol{v}})(-{\boldsymbol{p}}-{\boldsymbol{q}})\},&\mbox{if }\langle-{\boldsymbol{p}}-{\boldsymbol{q}},(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|})\rangle\geq 0,\\ \{{\boldsymbol{0}},B({\boldsymbol{v}})(-{\boldsymbol{p}}-{\boldsymbol{q}})\},&\mbox{ otherwise},\end{cases}\quad\forall\,{\boldsymbol{p}},{\boldsymbol{q}}\in{\mathds{R}}^{1+r},$ (3.7)
where the definition of $B({\boldsymbol{v}})$ comes from (3.2). Note that (3.7) holds if and only if
${\boldsymbol{q}}=\frac{\varrho}{2}\begin{pmatrix}1\\ \frac{\bar{{\boldsymbol{v}}}}{\|\bar{{\boldsymbol{v}}}\|}\end{pmatrix}\begin{pmatrix}1,\frac{\bar{{\boldsymbol{v}}}^{\top}}{\|\bar{{\boldsymbol{v}}}\|}\end{pmatrix}{({\boldsymbol{p}}+{\boldsymbol{q}})}\quad\mbox{with}\quad\varrho\in\begin{cases}[0,1],&\mbox{if }-\langle{\boldsymbol{p}},(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|})\rangle\geq\langle{\boldsymbol{q}},(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|})\rangle,\\ \{0,1\},&\mbox{ otherwise}.\end{cases}$ (3.8)
@@
$\begin{array}[]{ll}(1-\varrho)\langle(1;\frac{\bar{{\boldsymbol{v}}}}{\|\bar{{\boldsymbol{v}}}\|}),{\boldsymbol{q}}\rangle=\varrho\langle(1;\frac{\bar{{\boldsymbol{v}}}}{\|\bar{{\boldsymbol{v}}}\|}),{{\boldsymbol{p}}}\rangle.\end{array}$ (3.9)
Consequently, one can equivalently reformulate (3.8) to the condition that
@@
$(1-\varrho)\tau\begin{pmatrix}1\\ \frac{\bar{{\boldsymbol{v}}}}{\|\bar{{\boldsymbol{v}}}\|}\end{pmatrix}=\frac{\varrho}{2}\begin{pmatrix}1\\ \frac{\bar{{\boldsymbol{v}}}}{\|\bar{{\boldsymbol{v}}}\|}\end{pmatrix}\begin{pmatrix}1,\frac{\bar{{\boldsymbol{v}}}^{\top}}{\|\bar{{\boldsymbol{v}}}\|}\end{pmatrix}{{\boldsymbol{p}}}\quad\mbox{with}\quad\varrho\in\begin{cases}[0,1],&\mbox{if }-\langle{\boldsymbol{p}},(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|})\rangle\geq 2\tau,\\ \{0,1\},&\mbox{ otherwise}.\end{cases}$ (3.10)
@@
    If $\tau=0$, one can take $\varrho=0$ such that (3.10) holds for any ${\boldsymbol{p}}\in{\mathds{R}}^{1+r}$. Thus, ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{0}})={\mathds{R}}^{1+r}$.
@@
    If $\tau>0$, one can not take $\varrho=0$, or else (3.9) fails to hold. When $\varrho\in(0,1)$, by (3.9) one should have $\langle(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|}),{\boldsymbol{p}}\rangle>0$, but one can not find a vector ${\boldsymbol{p}}\in{\mathds{R}}^{1+r}$ such that $-\langle{\boldsymbol{p}},(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|})\rangle\geq 2\tau$. Therefore, one can only take $\varrho=1$, so that $\langle(1;\frac{\bar{{\boldsymbol{v}}}}{\|\bar{{\boldsymbol{v}}}\|}),{{\boldsymbol{p}}}\rangle=0$ by (3.9). In this case, it is easy to see that (3.10) holds. Thus, ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})=\{{\boldsymbol{p}}\mid\langle{\boldsymbol{p}},{\boldsymbol{q}}\rangle=0\}$.
@@
    If $\tau<0$, one has $\varrho\neq 0$ by the equality in (3.10). When $\varrho=1$, (3.10) holds if and only if $\langle(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|}),{\boldsymbol{p}}\rangle=0$. When $\varrho\in(0,1)$, one has $-\langle(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|}),{\boldsymbol{p}}\rangle=-\frac{1-\varrho}{\varrho}\langle(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|}),{\boldsymbol{q}}\rangle=\frac{-2\tau(1-\varrho)}{\varrho}>2\tau$ if (3.9) holds. Thus, (3.10) holds if and only if there exists $\varrho\in(0,1)$ such that $\frac{\varrho}{2(1-\varrho)\tau}\begin{pmatrix}1,\frac{\bar{{\boldsymbol{v}}}^{\top}}{\|\bar{{\boldsymbol{v}}}\|}\end{pmatrix}{{\boldsymbol{p}}}=1$. Since $\tau<0$, we know that (3.10) holds if and only if $\langle(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|}),{\boldsymbol{p}}\rangle<0$. Consequently, one has in this case that ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})=\{{\boldsymbol{p}}\mid\langle{\boldsymbol{p}},{\boldsymbol{q}}\rangle\geq 0\}$.
@@
(6a) Since ${\boldsymbol{u}}={\boldsymbol{v}}={\boldsymbol{0}}$, by (3.1) and Part $(6)$ of Lemma 3.1 one has ${\boldsymbol{0}}\in{\partial}_{B}\Pi_{{\mathcal{Q}}}({\boldsymbol{u}}+{\boldsymbol{v}}){\boldsymbol{r}}\subseteq{\mathcal{D}}^{*}\Pi_{{\mathcal{Q}}}({\boldsymbol{u}}+{\boldsymbol{v}})({\boldsymbol{r}})$ for all ${\boldsymbol{r}}\in{\mathds{R}}^{1+r}$. Then by Lemma 2.3 one has ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{0}})={\mathds{R}}^{1+r}$.
@@
${\boldsymbol{s}}-{\boldsymbol{r}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})(-{\boldsymbol{s}}),\quad\forall\,{\boldsymbol{r}}\in{\mathds{R}}^{1+r},$ (3.11)

### v3

@@ 3 Coderivatives of the second-order cone
(13) $\partial_{B}\Pi_{{\mathcal{Q}}}({\boldsymbol{0}})=\{I_{1+r},O_{1+r}\}\cup\left\{\frac{1}{2}\begin{pmatrix}1&{\boldsymbol{w}}^{\top}\\ {\boldsymbol{w}}&2\alpha I_{r}+(1-2\alpha){\boldsymbol{w}}{\boldsymbol{w}}^{\top}\end{pmatrix}\,\middle|\,\begin{array}[]{ll}{\boldsymbol{w}}\in{\mathds{R}}^{r},\|{\boldsymbol{w}}\|=1,\\ \alpha\in[0,1]\end{array}\right\}.$
@@
(14) $A({\boldsymbol{z}}):=I_{1+r}-\frac{1}{2}\begin{pmatrix}-1\\ \frac{\bar{{\boldsymbol{z}}}}{\|\bar{{\boldsymbol{z}}}\|}\end{pmatrix}\begin{pmatrix}-1&\frac{\bar{{\boldsymbol{z}}}^{\top}}{\|\bar{{\boldsymbol{z}}}\|}\end{pmatrix}\quad\mbox{and}\quad B({\boldsymbol{z}}):=\frac{1}{2}\begin{pmatrix}1\\ \frac{\bar{{\boldsymbol{z}}}}{\|\bar{{\boldsymbol{z}}}\|}\end{pmatrix}\begin{pmatrix}1&\frac{\bar{{\boldsymbol{z}}}^{\top}}{\|\bar{{\boldsymbol{z}}}\|}\end{pmatrix}.$
@@
(15) ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})=\{{\boldsymbol{p}}\mid{\boldsymbol{p}}=-{\boldsymbol{q}}-{\boldsymbol{r}},-W{\boldsymbol{r}}={\boldsymbol{q}}\}=\{{\boldsymbol{p}}\mid W{\boldsymbol{p}}=(I-W){\boldsymbol{q}}\},\quad\forall\,{\boldsymbol{q}}\in{\mathds{R}}^{1+r}.$
@@
which, together with (15), implies the explicit formulation of ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})$.
@@
(16) $-{\boldsymbol{q}}\in\begin{cases}{\sf conv}\{-{\boldsymbol{p}}-{\boldsymbol{q}},A({\boldsymbol{u}})(-{\boldsymbol{p}}-{\boldsymbol{q}})\},&\mbox{if }\langle(-{\boldsymbol{p}}-{\boldsymbol{q}}),(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|})\rangle\geq 0,\\ \{-{\boldsymbol{p}}-{\boldsymbol{q}},A({\boldsymbol{u}})(-{\boldsymbol{p}}-{\boldsymbol{q}})\},&\mbox{ otherwise},\end{cases}\quad\forall\,{\boldsymbol{p}},{\boldsymbol{q}}\in{\mathds{R}}^{1+r},$
@@
where the definition of $A({\boldsymbol{u}})$ comes from (14). Note that (16) holds if and only if
@@
(17) ${\boldsymbol{p}}=\frac{\varrho}{2}\begin{pmatrix}1\\ -\frac{\bar{{\boldsymbol{u}}}}{\|\bar{{\boldsymbol{u}}}\|}\end{pmatrix}\begin{pmatrix}1,-\frac{\bar{{\boldsymbol{u}}}^{\top}}{\|\bar{{\boldsymbol{u}}}\|}\end{pmatrix}{({\boldsymbol{p}}+{\boldsymbol{q}})},\ \varrho\in\begin{cases}[0,1],&\mbox{if }-\langle{\boldsymbol{p}},(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|})\rangle\geq\langle{\boldsymbol{q}},(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|})\rangle,\\ \{0,1\},&\mbox{ otherwise}.\end{cases}$
Moreover, (17) implies
(18) $\begin{array}[]{ll}(1-\varrho)\langle(1;-\frac{\bar{{\boldsymbol{u}}}}{\|\bar{{\boldsymbol{u}}}\|}),{\boldsymbol{p}}\rangle=\varrho\langle(1;-\frac{\bar{{\boldsymbol{u}}}}{\|\bar{{\boldsymbol{u}}}\|}),{{\boldsymbol{q}}}\rangle.\end{array}$
@@
For a given vector ${\boldsymbol{q}}\in{\mathds{R}}^{1+r}$, we consider the following three cases, based on the value of $\langle(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}),{\boldsymbol{q}}\rangle$, to get the equivalent conditions of ${\boldsymbol{p}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})$ from (17):
@@
    If $\langle(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}),{\boldsymbol{q}}\rangle>0$, one can not take $\varrho=1$, or else (18) fails to hold. When $\varrho\in(0,1)$, by (18) one should have $\langle(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}),{\boldsymbol{p}}\rangle>0$, but one can not find a vector ${\boldsymbol{p}}$ such that (17) holds in this case. Therefore, one can only take $\varrho=0$, so that ${\boldsymbol{p}}={\boldsymbol{0}}$ by the equality in (17). In this case, it is easy to see that (17) holds. Thus, ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})=\{{\boldsymbol{0}}\}$.
@@
    If $\langle(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}),{\boldsymbol{q}}\rangle=0$, by (18) one has $\langle(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}),{\boldsymbol{p}}\rangle=0$ whenever $\varrho\in[0,1)$. In this case, (17) holds if and only if ${\boldsymbol{p}}={\boldsymbol{0}}$. Moreover, when $\varrho=1$, (18) always holds. In this case, (17) holds if and only if ${\boldsymbol{p}}=\frac{1}{2}(1,-\frac{\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|}){\boldsymbol{p}}(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|})$. Thus, ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})=\{\tau(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|})\mid\tau\in{\mathds{R}}\}$.
@@
    If $\langle(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}),{\boldsymbol{q}}\rangle<0$, one can not take $\varrho=1$, or else (18) fails to hold. When $\varrho=0$, it is easy to see that (17) holds if and only if ${\boldsymbol{p}}={\boldsymbol{0}}$. When $\varrho\in(0,1)$, by (18) one should have $\langle(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}),{\boldsymbol{p}}\rangle<0$. Therefore, (17) holds if and only if ${\boldsymbol{p}}=\tau(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|})$ with $\tau:=\frac{\varrho}{2}(1,-\frac{\bar{\boldsymbol{u}}^{\top}}{\|\bar{\boldsymbol{u}}\|})({\boldsymbol{p}}+{\boldsymbol{q}})\leq 0$. Note that such a condition is valid for any $\tau<0$ since $\varrho=\frac{2\tau}{2\tau+\langle(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|}),{\boldsymbol{q}}\rangle}\in(0,1)$ always holds. Thus, ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})=\{\tau(1;-\frac{\bar{\boldsymbol{u}}}{\|\bar{\boldsymbol{u}}\|})\mid\tau\leq 0\}$.
@@
(19) $-{\boldsymbol{q}}\in\begin{cases}{\sf conv}\{{\boldsymbol{0}},B({\boldsymbol{v}})(-{\boldsymbol{p}}-{\boldsymbol{q}})\},&\mbox{if }\langle-{\boldsymbol{p}}-{\boldsymbol{q}},(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|})\rangle\geq 0,\\ \{{\boldsymbol{0}},B({\boldsymbol{v}})(-{\boldsymbol{p}}-{\boldsymbol{q}})\},&\mbox{ otherwise},\end{cases}\quad\forall\,{\boldsymbol{p}},{\boldsymbol{q}}\in{\mathds{R}}^{1+r},$
where the definition of $B({\boldsymbol{v}})$ comes from (14). Note that (19) holds if and only if
(20) ${\boldsymbol{q}}=\frac{\varrho}{2}\begin{pmatrix}1\\ \frac{\bar{{\boldsymbol{v}}}}{\|\bar{{\boldsymbol{v}}}\|}\end{pmatrix}\begin{pmatrix}1,\frac{\bar{{\boldsymbol{v}}}^{\top}}{\|\bar{{\boldsymbol{v}}}\|}\end{pmatrix}{({\boldsymbol{p}}+{\boldsymbol{q}})}\quad\mbox{with}\quad\varrho\in\begin{cases}[0,1],&\mbox{if }-\langle{\boldsymbol{p}},(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|})\rangle\geq\langle{\boldsymbol{q}},(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|})\rangle,\\ \{0,1\},&\mbox{ otherwise}.\end{cases}$
@@
(21) $\begin{array}[]{ll}(1-\varrho)\langle(1;\frac{\bar{{\boldsymbol{v}}}}{\|\bar{{\boldsymbol{v}}}\|}),{\boldsymbol{q}}\rangle=\varrho\langle(1;\frac{\bar{{\boldsymbol{v}}}}{\|\bar{{\boldsymbol{v}}}\|}),{{\boldsymbol{p}}}\rangle.\end{array}$
Consequently, one can equivalently reformulate (20) to the condition that
@@
(22) $(1-\varrho)\tau\begin{pmatrix}1\\ \frac{\bar{{\boldsymbol{v}}}}{\|\bar{{\boldsymbol{v}}}\|}\end{pmatrix}=\frac{\varrho}{2}\begin{pmatrix}1\\ \frac{\bar{{\boldsymbol{v}}}}{\|\bar{{\boldsymbol{v}}}\|}\end{pmatrix}\begin{pmatrix}1,\frac{\bar{{\boldsymbol{v}}}^{\top}}{\|\bar{{\boldsymbol{v}}}\|}\end{pmatrix}{{\boldsymbol{p}}}\quad\mbox{with}\quad\varrho\in\begin{cases}[0,1],&\mbox{if }-\langle{\boldsymbol{p}},(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|})\rangle\geq 2\tau,\\ \{0,1\},&\mbox{ otherwise}.\end{cases}$
@@
    If $\tau=0$, one can take $\varrho=0$ such that (22) holds for any ${\boldsymbol{p}}\in{\mathds{R}}^{1+r}$. Thus, ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{0}})={\mathds{R}}^{1+r}$.
@@
    If $\tau>0$, one can not take $\varrho=0$, or else (21) fails to hold. When $\varrho\in(0,1)$, by (21) one should have $\langle(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|}),{\boldsymbol{p}}\rangle>0$, but one can not find a vector ${\boldsymbol{p}}\in{\mathds{R}}^{1+r}$ such that $-\langle{\boldsymbol{p}},(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|})\rangle\geq 2\tau$. Therefore, one can only take $\varrho=1$, so that $\langle(1;\frac{\bar{{\boldsymbol{v}}}}{\|\bar{{\boldsymbol{v}}}\|}),{{\boldsymbol{p}}}\rangle=0$ by (21). In this case, it is easy to see that (22) holds. Thus, ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})=\{{\boldsymbol{p}}\mid\langle{\boldsymbol{p}},{\boldsymbol{q}}\rangle=0\}$.
@@
    If $\tau<0$, one has $\varrho\neq 0$ by the equality in (22). When $\varrho=1$, (22) holds if and only if $\langle(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|}),{\boldsymbol{p}}\rangle=0$. When $\varrho\in(0,1)$, one has $-\langle(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|}),{\boldsymbol{p}}\rangle=-\frac{1-\varrho}{\varrho}\langle(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|}),{\boldsymbol{q}}\rangle=\frac{-2\tau(1-\varrho)}{\varrho}>2\tau$ if (21) holds. Thus, (22) holds if and only if there exists $\varrho\in(0,1)$ such that $\frac{\varrho}{2(1-\varrho)\tau}\begin{pmatrix}1,\frac{\bar{{\boldsymbol{v}}}^{\top}}{\|\bar{{\boldsymbol{v}}}\|}\end{pmatrix}{{\boldsymbol{p}}}=1$. Since $\tau<0$, we know that (22) holds if and only if $\langle(1;\frac{\bar{\boldsymbol{v}}}{\|\bar{\boldsymbol{v}}\|}),{\boldsymbol{p}}\rangle<0$. Consequently, one has in this case that ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{q}})=\{{\boldsymbol{p}}\mid\langle{\boldsymbol{p}},{\boldsymbol{q}}\rangle\geq 0\}$.
@@
(6a) Since ${\boldsymbol{u}}={\boldsymbol{v}}={\boldsymbol{0}}$, by (13) and Part $(6)$ of Lemma 3.1 one has ${\boldsymbol{0}}\in{\partial}_{B}\Pi_{{\mathcal{Q}}}({\boldsymbol{u}}+{\boldsymbol{v}}){\boldsymbol{r}}\subseteq{\mathcal{D}}^{*}\Pi_{{\mathcal{Q}}}({\boldsymbol{u}}+{\boldsymbol{v}})({\boldsymbol{r}})$ for all ${\boldsymbol{r}}\in{\mathds{R}}^{1+r}$. Then by Lemma 2.4 one has ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})({\boldsymbol{0}})={\mathds{R}}^{1+r}$.
@@
(23) ${\boldsymbol{s}}-{\boldsymbol{r}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}}({\boldsymbol{u}},{\boldsymbol{v}})(-{\boldsymbol{s}}),\quad\forall\,{\boldsymbol{r}}\in{\mathds{R}}^{1+r},$